# Are penalty shootouts better than a coin toss? Evidence from European football

### v2

@@ Abstract
Penalty shootouts play an important role in the knockout stage of major football tournaments, especially since the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule in its club competitions. Inspired by this rule change, our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, stronger teams, defined first by Elo ratings, do not perform better than their weaker opponents. Consequently, penalty shootouts are equivalent to a perfect lottery in top European football.
The outcome of penalty shoot-outs is often referred to as a “lottery”, suggesting that luck, rather than the skill level of the player, predetermines outcome success.111 Source: Wood et al., (2015, Abstract).
@@ 1 Introduction
The penalty shootout is the ultimate tie-breaking rule in (association) football and several other sports to decide which team qualifies in a knockout match. Penalty shootouts usually take place if the result is tied in both regular and extra time. Penalty shootouts directly follow regular time in some tournaments, too. For instance, in the 2025 Leagues Cup, games went straight to a penalty shootout after a tie, and teams were awarded three points for a win in regular time, two points for a win by a penalty shootout, and one point for a loss by a penalty shootout (Leagues Cup,, 2025).
@@
All of these issues have partially been investigated in the existing literature, as revealed by Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, psychological momentum enjoyed by the comeback team is identified in a novel way, according to which team scored the last goal. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose. But the first approach is inappropriate in international competitions, while betting odds suffer from several biases (Braun and Kvasnicka,, 2013; Cain et al.,, 2000; Feddersen et al.,, 2017; Winkelmann et al.,, 2021), and may be less persuasive for decision-makers than a measure given by an “objective” mathematical formula.
@@
Our findings could be especially important for the Union of European Football Associations (UEFA) when they decide on rule changes in the future. In 2021, UEFA abolished the so-called away goals rule, which awarded the prize (qualification for the next round) in a two-legged match to the team that scored more away goals. This change has increased the probability of reaching a penalty shootout by approximately ten percentage points, from below 5% to nearly 15% (Forrest et al.,, 2025). In addition, UEFA has recently given serious consideration to removing extra time in order to ease the burden of the best players (Ames,, 2025). Since a penalty shootout is equivalent to a perfect lottery in UEFA club competitions, such a measure would favour weaker teams and teams playing away, level the playing field, and decrease the dominance of the top clubs.
The paper is structured as follows. An overview of related studies is given in Section 2. Materials and methods are presented in Section 3. Section 4 reveals and discusses the results, while Section 5 concludes.
@@ 2 Related literature
Regarding the existence of a penalty-specific home advantage, Apesteguia and Palacios-Huerta, (2010) show no difference in the chances of home and away teams for 129 shootouts between 1976 and 2003. Kocher et al., (2012) reinforce this based on 540 shootouts between 1970 and 2003. Analogously, playing at home does not influence the probability of winning in 252 shootouts from French cup competitions (Arrondel et al.,, 2019). The results of Wunderlich et al., (2020) suggest no home advantage in 1067 penalty shootouts that took place between 2004 and 2018. The logistic regression of Bahamonde-Birke and Bahamonde-Birke, (2023) based on a sample of 471 shootouts demonstrates the absence of such an effect, too.
@@
Brinkschulte et al., (2023) examine the effect of skill on the success of individual penalty kicks. Based on 1574 penalties in major international tournaments, the probability that the goalkeeper saves a penalty that is shot on target decreases if a highly skilled player takes the kick. Skill is a composite measure of player market value, goalkeeper market value, and team quality derived from betting odds.
@@ 3 Data and methodology
The connection between the relative strength of teams and the probability of winning the penalty shootout is assessed by logistic regressions, too. The binary dependent variable is the home win/loss. Denote the Elo rating of the home and the away team by $E_{i}$ and $E_{j}$, respectively. The only independent variable $\Delta_{ij}$ is either the difference of Elo ratings ($\Delta_{ij}^{(1)}=E_{i}-E_{j}$), or the winning probability implied by the Elo ratings according to the Elo equation of Football Club Elo Ratings (http://clubelo.com/system):
$\Delta_{ij}^{(2)}=\frac{1}{1+10^{-(E_{i}-E_{j})/400}}.$
Note that $\Delta_{ji}^{(1)}=-\Delta_{ij}^{(1)}$ and $\Delta_{ji}^{(2)}=1-\Delta_{ij}^{(2)}$.
@@
$P_{h}=\frac{1}{1+e^{\beta_{0}+\beta_{1}\Delta}},$
where $P_{h}$ is the probability that the home team wins the penalty shootout, and $\Delta$ is the strength difference of the opposing teams. Matches played at a neutral venue are considered according to the designated home team.
a Difference in Elo ratings $\Delta^{(1)}$
b Difference in winning probabilities $\Delta^{(2)}$
0–0.050.05–0.10.1–0.150.15–0.20.2–0.250.25–0.30.3–0.350.35–0.40.4–0.450.45–0.50.5–0.550.55–0.60.6–0.650.65–0.70.7–0.750.75–0.80.8–0.850.85–0.90.9–0.950.95–1010102020Home team winning probability minus away team winning probabilityNumber of penalty shootouts
Figure 2: Distribution of the difference between team strengths
Figure 2 shows histograms of strength difference $\Delta$: Figure 2.a if the difference is measured directly by Elo ratings ($\Delta^{(1)}$), and Figure 2.b if it is based on the implied winning probabilities ($\Delta^{(2)}$). Although all matches are decided by a penalty shootout, the database is quite heterogeneous, the winning probability of the home team lies between 40% and 60% in less than 30% of the sample.
@@ 4 Results and discussion
First-mover winHome win (extended)Home win (restricted)Comeback win (all)Comeback win (last)40404545505055556060Probability (%)
@@
Elo difference >0>0Elo difference >25>25Elo difference >50>50Elo difference >75>75Elo difference >100>10040404545505055556060Probability (%)
@@
Figure 4 focuses on the winning probability of the stronger team based on Football Club Elo Ratings. Five definitions of the favourite are used by gradually restricting the sample according to the minimal difference in the Elo ratings of the two teams. However, even teams that have at least 100 points more than their opponents fail to win 50% of their penalty shootouts. These calculations can be reproduced on the basis of Table A.2 in Appendix A.2, which also reports the associated $p$-values.
@@
As we have seen in Section 3, the number of penalty shootouts has substantially increased from the 2020/21 season. Thus, all binomial tests have also been conducted for the first 20 and the last six seasons, respectively. According to Tables A.3A.6 in the Appendix, the conclusions remain unchanged if we focus on the beginning or the end of the dataset. The only mildly significant value appears for home advantage based on the extended dataset: home teams are less likely to win a penalty shootout than their opponents between 2020 and 2025 if the matches played behind closed doors are taken into account (Table A.5). In these 20 one-legged matches, only seven was won by the home team, which is significantly below 50% at a significance level of 10%. This can be explained by choking under pressure (Harb-Wu and Krumer,, 2019), which was probably much stronger in these matches than usual, despite the absence of spectators, as, in contrast to the standard rule, the randomly drawn home team should not have played away. Even this marginally significant effect disappears when only proper two-legged matches are considered under the restricted definition of home venue (Table A.5).
Table 2: Logistic regression models: the effect of playing at home and team strength
Strength difference $\Delta$ Elo rating $\Delta^{(1)}$ Winning probability $\Delta^{(2)}$ Coefficient -0.121 -0.027 $\beta_{0}$ Standard error 0.123 0.307 $p$-value 0.326 0.930 Coefficient -0.004 -0.194 $\beta_{1}$ Standard error 0.069 0.582 $p$-value 0.959 0.739 Number of observations 266 266 In the model Elo rating, strength difference $\Delta^{(1)}$ is the Elo rating of the home team minus the Elo rating of the away team divided by 100. In the model Elo rating, strength difference $\Delta^{(2)}$ is the winning probability of the home team based on the Elo ratings of the two teams. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
The lack of positive impact is in stark contrast to the findings of the previous literature: all existing papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025) demonstrate a significantly higher winning probability for a “sufficiently” stronger team. Thus, Table 2 investigates the same issue by logistic regressions similar to Wunderlich et al., (2020). Both coefficients are insignificant, that is, neither the field of the match nor the strength of the teams influence success in penalty shootouts, independent of how the strength difference is measured. Furthermore, the estimations of $\beta_{0}$ and $\beta_{1}$ are negative in accordance with Figures 3 and 4, making the implications regarding home advantage and team strength even more robust.
@@
200020002002200220042004200620062008200820102010201220122014201420162016201820182020202020222022202420240.010.010.10.111Starting year of the first season in the estimationpp-valuepp-value of the coefficient $\beta_{0}$ for the constant   $p$-value of the coefficient $\beta_{1}$ for $\Delta$
@@
200020002002200220042004200620062008200820102010201220122014201420162016201820182020202020222022202420240.010.010.10.111Starting year of the first season in the estimationpp-valuepp-value of the coefficient $\beta_{0}$ for the constant   $p$-value of the coefficient $\beta_{1}$ for $\Delta$
@@
Notes: The dotted (dashed) line shows significance at 5% (1%) level.
Seasons are identified by their starting year.
Last but not least, we have assessed whether team strength could have a significant effect if the sample is restricted to the recent seasons. Hence, 26 logistic regressions are estimated in both versions such that the sample begins in the 20xx/20xx+1 season and ends in 2025. Figure 5 plots the corresponding $p$-values for the coefficients $\beta_{0}$ and $\beta_{1}$. The winning probability becomes different from a coin toss only if the sample consists of the last season 2025/26—when the number of observations is 14, and a higher Elo rating seems to be disadvantageous. Consequently, in contrast to domestic cups, stronger teams are not expected to win more penalty shootouts in UEFA club competitions.

### v3

# Are penalty shootouts better than a coin toss? Evidence from international club football in Europe

@@ Abstract
Penalty shootouts play an important role in the knockout stage of major football tournaments, especially since the 2021/22 season, when the Union of European Football Associations (UEFA) scrapped the away goals rule in its club competitions. Inspired by this rule change, our paper examines whether the outcome of a penalty shootout can be predicted in UEFA club competitions. Based on all shootouts between 2000 and 2025, we find no evidence for the effect of the kicking order, the field of the match, or psychological momentum. In contrast to previous results, stronger teams, defined first by Elo ratings, do not perform better than their weaker opponents. Consequently, penalty shootouts seem to be close to a coin toss in top European club football.
@@ 1 Introduction
The penalty shootout is the ultimate tie-breaking rule in (association) football and several other sports to decide which team qualifies in a knockout match. Penalty shootouts usually take place if the result is tied in both regular and extra time. Penalty shootouts can directly follow regular time, too: in the 2025 Leagues Cup, teams were awarded three points for a win in regular time, two points for a win by a penalty shootout, and one point for a loss by a penalty shootout (Leagues Cup,, 2025).
@@
All of these issues have partially been investigated in the existing literature, as revealed by Section 2. However, our research has three innovative aspects. First, we analyse a relatively homogeneous sample provided by matches played in UEFA club competitions without any matches from national cups, where the strengths of the opponents can vary widely. Second, psychological momentum enjoyed by the comeback team is identified in a novel way, according to which team scored the last goal. Third, team strength is measured by Football Club Elo Ratings, a well-established statistical method to create ratings based on past performance (Aldous,, 2017; Gomes de Pinho Zanco et al.,, 2024; van Eetvelde and Ley,, 2019). Previous studies used either the division of the team (Arrondel et al.,, 2019; Krumer,, 2020), or betting odds (Wunderlich et al.,, 2020; Pipke,, 2025) for this purpose.
@@
Our findings could be especially important for the Union of European Football Associations (UEFA) when they decide on rule changes in the future. In 2021, UEFA abolished the so-called away goals rule, which awarded the prize (qualification for the next round) in a two-legged match to the team that scored more away goals. This change has increased the probability of reaching a penalty shootout by approximately ten percentage points, from below 5% to nearly 15% (Forrest et al.,, 2025). In addition, UEFA has recently given serious consideration to removing extra time in order to ease the burden of the best players (Ames,, 2025). Since the outcome of a penalty shootout approximates a coin toss in UEFA club competitions with respect to these observables, such a measure would favour weaker teams and teams playing away, level the playing field, and decrease the dominance of the top clubs.
The paper is structured as follows. An overview of related studies is given in Section 2. The data and the methodology are described in Sections 3 and 4, respectively. Section 5 presents and discusses the results, while Section 6 concludes.
@@ 2 Related literature
Regarding the existence of a penalty-specific home advantage, Apesteguia and Palacios-Huerta, (2010) show no difference in the chances of home and away teams for 129 shootouts between 1976 and 2003. Kocher et al., (2012) reinforce this based on 540 shootouts between 1970 and 2003. Analogously, playing at home does not influence the probability of winning in 252 shootouts from French cup competitions (Arrondel et al.,, 2019). The results of Wunderlich et al., (2020) suggest no home advantage in 1067 penalty shootouts that took place between 2004 and 2018. The logistic regression of Bahamonde-Birke and Bahamonde-Birke, (2023) based on a sample of 471 shootouts demonstrates the absence of such an effect, too. Even though van Ours and van Tuijl, (2024) state that a penalty shootout does not seem to be a lottery in the Dutch domestic cup, the 20 home wins out of the 33 shootouts (61%) are not significantly different from 50%.
@@
Brinkschulte et al., (2023) examine the effect of skill on the success of individual penalty kicks. Skill is a composite measure of player market value, goalkeeper market value, and team quality derived from betting odds. Based on 1574 penalties in major international tournaments, the probability that the goalkeeper saves a penalty that is shot on target decreases if a highly skilled player takes the kick.
@@ 3 Data
It may be argued that betting odds could be a better predictor of team strength in penalty shootouts. However, betting odds are known to suffer from several biases (Braun and Kvasnicka,, 2013; Cain et al.,, 2000; Feddersen et al.,, 2017; Winkelmann et al.,, 2021). Furthermore, the betting odds published by bookmakers are usually related to the outcome of one particular match, but our database mostly contains the second legs of knockout ties. Therefore, the favourite team identified by betting odds might not be the stronger team because of home advantage. Even though Wunderlich et al., (2020, Section 2.3) suggests a method to adjust this bias, which allows the classification of a slight favourite as the slightly weaker team, such a procedure is not necessarily reliable in two-legged matches. The reason is the incentives of the teams: one team could be satisfied by a small loss if it still implies its qualification. For instance, if the away team won the first match by two goals, it does not have powerful incentives to attack in the second match even if it has scored one goal less than the other team. The betting odds of a home/away win certainly reflect the effect of these incentives, which have substantially changed from the 2021/22 season due to the abolition of the away goals rule.
Compared to Elo ratings, betting odds can indeed reflect better the current form of the teams by taking, for example, injuries, into account. However, it is far from clear that these factors are relevant to penalty shootouts, where success is based on individual skills, not on collaboration within a team. The use of Elo ratings, given by an “objective” mathematical formula, may also be more persuasive for decision-makers than betting odds driven by market sentiment. Finally, Football Club Elo Ratings have recently been demonstrated to outperform UEFA club coefficients, the official measure of team strength in these competitions, with respect to predictive accuracy (Csató,, 2024). Based on these arguments, we follow the approach of previous papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025), and do not consider alternative measures of team ability.
4 Methodology
@@
The connection between the relative strength of teams and the probability of winning the penalty shootout is assessed by logistic regressions, too. The binary dependent variable is the home win/loss. Therefore, matches played on a neutral venue are removed due to the small sample size, similar to (Wunderlich et al.,, 2020). Denote the Elo rating of the home and the away team by $E_{h}$ and $E_{a}$, respectively. The independent variable $\Delta_{ha}$ is either the difference of Elo ratings ($\Delta_{ha}^{(1)}=E_{h}-E_{a}$), or the winning probability of the home team implied by the Elo ratings according to the Elo equation of Football Club Elo Ratings (http://clubelo.com/system):
$\Delta_{ha}^{(2)}=\frac{1}{1+10^{-\left(E_{h}-E_{a}\right)/400}}.$
Note that $\Delta_{ah}^{(1)}=-\Delta_{ha}^{(1)}$ and $\Delta_{ah}^{(2)}=1-\Delta_{ha}^{(2)}$.
a Distribution of strength difference $\Delta^{(1)}$
b Distribution of strength difference $\Delta^{(2)}$
0–0.050.05–0.10.1–0.150.15–0.20.2–0.250.25–0.30.3–0.350.35–0.40.4–0.450.45–0.50.5–0.550.55–0.60.6–0.650.65–0.70.7–0.750.75–0.80.8–0.850.85–0.90.9–0.950.95–1010102020Winning probability of the home teamNumber of penalty shootouts
Figure 2: Distribution of the difference between team strengths
Figure 2 shows histograms of strength difference $\Delta$: Figure 2.a if the difference is measured directly by Elo ratings ($\Delta^{(1)}$), and Figure 2.b if it is based on the winning probability of the home team ($\Delta^{(2)}$). Although all matches are decided by a penalty shootout, the database is quite heterogeneous, the winning probability of the home team lies between 40% and 60% in less than 30% of the sample.
The first covariate is being the first-mover ($f_{h}=1$ if the home team kicks the first penalty). The second covariate is the comeback team ($c_{h}=1$ if the home team scores more goals or the last goal). However, including the comeback dummy reduces sample size since no team scores more goal if the final result is 0-0, and no team scores the last goal if the final result is 0-0. There is no need to control for match venue, the possible home (dis)advantage is reflected by the constant.
@@
$P_{h}=\frac{1}{1+e^{\beta_{0}+\beta_{1}\Delta+\beta_{2}f_{h}+\beta_{3}c_{h}}},$
where $P_{h}$ is the probability that the home team wins the penalty shootout, and $\Delta$ is the strength difference of the home and the away team. We consider 12 different regressions, depending on the definition of home venue (extended/restricted), the measure of strength difference (Elo rating/winning probability), and the momentum effect (ignored: $\beta_{3}=0$/more goals scored/last goal scored).
@@ 5 Results and discussion
Section 5.1 investigates potential success factors in penalty shootouts by binomial tests, that is, whether teams in a given category win significantly more/less often than 50% compared to teams in the other category. Section 5.2 considers more dimensions simultaneously via logistic regressions. In both cases, sensitivity analyses are also conducted. Finally, we provide some explanations for the results in Section 5.3.
5.1 Partial analysis
First-moverHome (extended)Home (restricted)Comeback (all)Comeback (last)40404545505055556060Winning probability (%)
@@ 5.1 Partial analysis
Elo difference >0>0Elo difference >25>25Elo difference >50>50Elo difference >75>75Elo difference >100>10040404545505055556060Winning probability (%)
@@
Figure 4 focuses on the winning probability of the stronger team based on Football Club Elo Ratings. Five definitions of the favourite are used by gradually restricting the sample according to the minimal difference in the Elo ratings of the two teams. However, even teams that have at least 100 points more than their opponents fail to win 50% of their penalty shootouts. These calculations can be reproduced on the basis of Table A.2 in Appendix A.2, which also reports the associated $p$-values. The lack of positive impact is in stark contrast to the findings of the previous literature: all existing papers (Arrondel et al.,, 2019; Krumer,, 2020; Wunderlich et al.,, 2020; Pipke,, 2025) demonstrate a significantly higher winning probability for a “sufficiently” stronger team.
@@
As we have seen in Section 3, the number of penalty shootouts has substantially increased from the 2020/21 season. Thus, all binomial tests have also been conducted for the first 20 and the last six seasons, respectively. According to Tables A.1A.2 in the Appendix, the conclusions remain unchanged if we focus on the beginning or the end of the dataset. The only significant value at the level of 5% (but not at the level of 1%) appears for home advantage based on the extended dataset: home teams are less likely to win a penalty shootout than their opponents between 2020 and 2025 if the matches played behind closed doors are taken into account (Table A.1). In these 20 one-legged matches, only seven were won by the home team, which is significantly below 50% at a significance level of 10%. This can be explained by choking under pressure (Harb-Wu and Krumer,, 2019), which was probably much stronger in these matches than usual, despite the absence of spectators, as, in contrast to the standard rule, the randomly drawn home team should not have played away. The marginally significant effect disappears when only proper two-legged matches are considered under the restricted definition of home venue (Table A.1). Tables A.3 and A.4 reinforce the insignificant impact of all factors considered by distinguishing the penalty shootouts played in the qualification and knockout phases.
Even though the binomial tests do not indicate a significant deviation from the theoretically expected 50%, this might be caused by a small sample size. Furthermore, a failure to reject the null hypothesis rules out only sufficiently large effects. Therefore, it is worth noting that a first-mover advantage can be observed to some extent, especially between 2000 and 2019, when the team kicking the first penalty had a higher probability of winning by more than 10 percentage points. On the other hand, while home teams won more penalty shootouts in the first 19 seasons, they might be disadvantaged recently: since 2020, the winning probability of the home team is lower by about 30% compared to the away team. Last but not least, a slight but consistent comeback team advantage is seen if the comeback team is identified by scoring the last goal, which is a novel definition in the literature. In particular, these teams enjoy an advantage of 9 (based on the whole sample) and 14 (in the last six seasons) percentage points over their opponents, albeit this is still statistically insignificant. Interestingly, the team kicking the first penalty or scoring the last goal is more likely to win—by about five percentage points—in matches played in the knockout stage.
5.2 Multi-dimensional analysis
Table 2: Logistic regression models for predicting success in penalty shootouts
a Extended definition of home-away matches
Measure of strength Elo difference $\Delta^{(1)}$ Home win probability $\Delta^{(2)}$
Constant ($\beta_{0}$) -0.269 -0.034 -0.501* -0.265 -0.065 -0.574
(0.179) (0.347) (0.250) (0.347) (0.586) (0.425)
Strength difference ($\beta_{1}$) 0.019 0.046 0.043 -0.009 0.095 0.154
(0.071) (0.101) (0.079) (0.595) (0.846) (0.659)
First-mover ($\beta_{2}$) 0.243 0.180 0.266 0.237 0.178 0.263
(0.253) (0.342) (0.275) (0.254) (0.342) (0.275)
Comeback (all goals, $\beta_{3}$) -0.122 -0.158
(0.397) (0.395)
Comeback (last goal, $\beta_{3}$) 0.402 0.387
(0.282) (0.281)
Observations 253 145 219 253 145 219
  • The dependent variable is 1 (0) if the home (away) team won the penalty shootout. Each column represents a separate regression.
  • Standard errors are in parentheses. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
  • Elo difference $\Delta^{(1)}$ is the Elo of the home team minus the Elo of the away team divided by 100.
b Restricted definition of home-away matches
Measure of strength Elo difference $\Delta^{(1)}$ Home win probability $\Delta^{(2)}$
Constant ($\beta_{0}$) -0.141 -0.034 -0.407 -0.080 -0.065 -0.399
(0.188) (0.347) (0.261) (0.372) (0.586) (0.455)
Strength difference ($\beta_{1}$) 0.005 0.046 0.025 -0.120 0.095 -0.001
(0.075) (0.101) (0.083) (0.628) (0.846) (0.695)
First-mover ($\beta_{2}$) 0.075 0.180 0.196 0.066 0.178 0.189
(0.265) (0.342) (0.285) (0.265) (0.342) (0.285)
Comeback (all goals, $\beta_{3}$) -0.122 -0.158
(0.397) (0.395)
Comeback (last goal, $\beta_{3}$) 0.325 0.309
(0.292) (0.291)
Observations 232 145 204 232 145 204
  • The dependent variable is 1 (0) if the home (away) team won the penalty shootout. Each column represents a separate regression.
  • Standard errors are in parentheses. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
  • Elo difference $\Delta^{(1)}$ is the Elo of the home team minus the Elo of the away team divided by 100.
Table 2 investigates the effect of strength difference between the teams on the probability of winning a penalty shootout by logistic regressions, similar to Wunderlich et al., (2020). We study only matches with a home and an away team (Table 2.a uses an extended and Table 2.b uses a restricted definition), and control for first-mover advantage. The consideration of the momentum effect reduces the number of observations since drawn matches or matches with a final result of 0-0 are excluded. Note that the model containing comeback teams identified by scoring more goals does not depend on the definition of home venue, as the 20 matches affected by the Covid-19 pandemic and a particular final are all played in one leg (see Section 3), thus, their final result should be a draw if they are finished by a penalty shootout.
In Table 2, only one significant coefficient can be found. In particular, the constant is negative if the extended definition of home venue is used and the comeback team is the team scoring the last goal. This is in line with Table A.1, which also indicates a home disadvantage since 2020 based on the extended set of matches. However, the $p$-value remains above 0.1 if the restricted definition of home venue is applied, showing the sensitivity of the finding to the definition of home venue. Albeit insignificant, the coefficient $\beta_{1}$ is always positive, while $\beta_{3}$ is negative/positive when the momentum effect is identified by scoring all goals/the last goal, as expected from the binomial tests in Figure 2. Crucially, the coefficient $\beta_{2}$ of strength difference is robustly insignificant.
@@ 5.2 Multi-dimensional analysis
200020002002200220042004200620062008200820102010201220122014201420162016201820182020202020222022202420240.010.010.10.111Starting year of the first season in the estimationpp-valuepp-value of the constant   $p$-value of $\beta_{1}$ for $\Delta^{(1)}\qquadpp$-value of $\beta_{2}$ for first-mover
@@
200020002002200220042004200620062008200820102010201220122014201420162016201820182020202020222022202420240.010.010.10.111Starting year of the first season in the estimationpp-valuepp-value of the constant   $p$-value of $\beta_{1}$ for $\Delta^{(2)}\qquadpp$-value of $\beta_{2}$ for first-mover
@@
Note: The dotted (dashed) line shows significance at 5% (1%) level.
Last but not least, we have assessed whether team strength could have a significant effect by focusing on the recent seasons. Hence, 25-25 logistic regressions are estimated by controlling for strength difference and first-mover effect ($\beta_{3}=0$ is assumed in order to maximise the number of observations): the sample begins in the 20xx/20xx+1 season and ends in 2025. Figure 5 plots the corresponding $p$-values for the coefficients $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. The winning probability becomes different from a coin toss when all seasons before 2023/24 are ignored—but then the sample size is not more than 64. Furthermore, $\beta_{1}$ is significant only if the last one or two season(s) are considered. Even then, both $\beta_{1}$ and $\beta_{2}$ are negative, which indicates rather a statistical anomaly caused by inadequate sample size. Consequently, in contrast to domestic cups, stronger teams are not expected to win more penalty shootouts in UEFA club competitions.
5.3 Discussion
At first sight, our findings may seem to contradict results from the previous literature. The detailed comparison of the estimations by Apesteguia and Palacios-Huerta, (2010) and Kocher et al., (2012) uncovers that the discrepancy is often caused by sampling issues, see Kocher et al., (2012). The majority of our sample consists of matches played in the last six seasons, which can explain at least some differences.
Regarding first-mover advantage, the results are fully in line with the conclusions of recent studies on large samples (Pipke,, 2025; Vollmer et al.,, 2024). Similar to us, no previous work has found evidence for home advantage. Even though Krumer, (2020) reports a momentum effect for away (but not for home) teams, there are two important differences between the two analyses. First, Krumer, (2020) measures momentum by the goal difference between the teams in regular time of the second leg; thus, it is not a binary variable as in our case. Second, the sample of Krumer, (2020) contains matches played between the 1970/71 and 2017/18 seasons, and there is powerful evidence for a decreasing home advantage in club football over time (Baker and McHale,, 2015). Therefore, away teams were more disadvantaged in matches played decades ago, hence, it is logical to assume that they enjoyed a stronger momentum effect if they managed to score more goals against their opponents.
In contrast to all existing studies, stronger teams are found not to win more often, which can be deemed counterintuitive. However, we do not see such a problem after a more detailed analysis. First, Arrondel et al., (2019) and Krumer, (2020) find that teams from a higher-ranked division have a higher probability of winning in a penalty shootout. Our sample is more homogeneous from this perspective: teams from the highest-ranked European leagues do not play in the UEFA Champions League qualification since the 2018/19 season, and at most one team participates in the UEFA Europa League and UEFA (Europa) Conference League qualification. Furthermore, the UEFA Champions League qualification was separated into two paths in the 2009/10 season such that champions from lower-ranked national associations are not allowed to play against non-champions from higher-ranked national associations (Csató,, 2022). Even though the knockout stage does not contain analogous constraints, the previous group stage/league phase can effectively remove weak clubs, guaranteeing that teams playing against each other are quite close in abilities. Second, although Wunderlich et al., (2020) report a significant impact for betting odds in domestic cups, its magnitude is relatively small, even extremely weaker teams have a probability of around 40% to win the penalty shootout. UEFA club competitions are more prestigious and lucrative than national clubs, thus strong teams probably field their best squad. Therefore, if the strength difference between the opposing teams is substantial, the match is less likely to be decided by a penalty shootout in UEFA club competitions than in national cups.
Any penalty shootout is clearly a situation with high stakes. In such an environment, choking under pressure can threaten the stronger team, which is expected to win. Indeed, professional biathlon athletes in the top quartile of the ability distribution miss significantly more shots when competing in their home country compared to competing abroad (Harb-Wu and Krumer,, 2019). In addition, task-specific practice could contribute to the success of penalty kicks (Navarro et al.,, 2013). This is especially important because, according to Jordet, (2009), players from countries with many international club titles or many internationally decorated players (such as England or Spain), spend less time preparing for their shots and, consequently, are less successful in kicking penalties than players from countries with a lower public status.